# Investigating Emotions, Social Roles, Well-being and Mobile Usage Behaviors During COVID-19 Home Confinement

### v1

@@ 3.1. Background
To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (54; 80). Our data collection was conducted during the second lockdown period with Stage 4 restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage 4 restrictions in metropolitan Melbourne (58).
Under the Stage 4 restrictions in force at the time our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (69; 77). This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviors of individuals during home confinement. The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
@@ 3.2. Participants
Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers. In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, std = 8.95). The age distribution of participants across genders can be found in Figure 2. It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample. While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behavior. Rather, we focus on individuals’ psychological states and smartphone usage behaviors during COVID-19 home confinement. Therefore, we believe that the potential oversample threat to the validity of our results is minimized. However, it is still important to note that our findings are specific to this particular population subset and may not be necessarily generalizable to other populations. Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
After the data collection, each participant was compensated a 25 AUD gift card as a token of appreciation for their time. Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage. Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants. Analysis of desktop usage behaviors will be included at the end of the discussion and explored further in future work.
@@ 3.3. Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (33) on their smartphones and desktops, respectively. These applications were specifically designed to record usage behaviors and enable background sensing and experience sampling, drawing inspiration from Van et al. (85). Both applications were optimized to minimize their consumption of computational power, energy, and storage resources. The applications utilized ESM, popping up a short questionnaire to participants every 90 minutes when their devices were actively in use between 7 am and 10 pm. This helps to obtain unbiased samples of participants’ activities, which served as the ground truth for our predictive models. Furthermore, an event-based ESM was triggered on the user’s smartphone after a minimum interaction of 10 minutes, with subsequent event-based ESMs occurring every 30 minutes.
@@ 3.3.1. Smartphone usage data
During the data collection, the background services of smartphones and desktops continuously tracked the following information: application usage (using Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (using Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (i.e. screen/battery status, ringer modes, location updates and physical activities). Specifically, the mobile app usage behavior was recorded every time the user opens the app. For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none). Table 1 displays an overview of the collected data for smartphone usage, ESM and end-of-day survey. In total, we have collected 502,485 recordings for smartphone usage data from 27 participants.
@@ 3.3.2. ESM responses
There are some popular questionnaires to measure well-being (39; 9), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (9) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users. Especially, the item ‘valence’ was rated with a 5-point Likert scale from ‘unhappy’ to ‘happy’, and the item ‘arousal’ was also rated with a 5-point Likert scale from ‘calm’ to ‘excited’. Both valence and arousal were used to represent the users’ emotions based on the circumplex model (74). Understanding the social role, interruptiblity and task help us understand how people manage their time and attention on the digital devices (3). The detailed items of the ESM can be found in Table 1 and the screenshot of the ESM on users’ smartphones was shown in Figure 3(a). Overall, we have collected 1,749 ESM responses from 29 participants. Of these, 1,522 responses were obtained from 25 participants using smartphones (2 participants did not contribute any ESM responses), while the remaining 227 responses were from 10 participants using desktops.
@@ 4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
In this section, we aim to answer RQ1: ”What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”. To achieve this, we analyze 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1. Furthermore, we explore the influence of COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
@@ 4.1. The Perceived Valence, Arousal and Social roles
The distribution of perceived valence, arousal, and social roles from the ESM responses is depicted in Figure 4. On average, participants reported a valence level of 3.46 (SD = 1.01) and an arousal level of 2.72 (SD = 1.11). Usually, the relatively high valance and low arousal refer to a state of calmness, relaxation and peace. This could be attributed to various factors during the COVID-19 home confinement. One possible reason is that people may feel relieved to be safe at home, and cherish the opportunity to spend more quality time with their family and loved ones, away from the potential risks of infection outside. Furthermore, the absence of the usual stressors of daily life, such as commuting, work pressures, and social obligations, may be contributing to a general sense of calm and well-being. Additionally, some people may have found comfort in the sense of community and solidarity that emerged during the pandemic, as people came together to support one another through difficult times. All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement. Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage 4 restrictions (strict home confinement) were lifted and steps for reopening commenced. Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges. For instance, prolonged periods of confinement may lead to feelings of boredom, loneliness, and frustration.
@@
Next, we investigate the distribution of social roles and their correlation with emotion during home confinement. A social role refers to a set of norms, expectations, rules, beliefs and behaviors that are associated with a particular social situation (70). These roles can significantly influence users’ behaviors and the type of information they prefer to receive (3). As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
Since social roles play a crucial role in shaping individuals’ well-being, we investigate the emotions of users in different social roles, as shown in Figure 4(c). We found that, when engaging in private-related roles, participants had the highest levels of valence and lowest levels of arousal compared to work-related roles and hybrid roles. This may be because private-related roles involve personal activities which can provide a sense of comfort and happiness, while work-related roles or hybrid roles may involve stress, deadlines, and pressure to perform. The results of Welch’s t-tests (94) further support these findings, indicating that there is a significant difference in valence levels between private-related roles and the other two roles (i.e., private vs. both, private vs. work), with p-value less than 0.05. The above findings provides insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (31).
@@ 4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
In total, 54.3%, 51.3% and 54.7% of end-of-day responses displayed that the lockdown has influenced their valence, energy, and work-life balance to some extent, as shown in Figure 5. For the influence of lockdown on valence, 13 participants reported that it had a negative effect, claiming they ‘felt quite anxious and sad’ as ‘staying at home has made me a little sad’. They complained that they were unable to see friends (9 responses), see families (5 responses), go outside (10 responses), and feel bored (6 responses), e.g., ‘I thought I might be able to see friends, but I did not’, ‘I want to go outside, to the park’, ‘I’m missing my family’, ‘It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped ’. Nevertheless, two felt the lockdown has a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
@@
Besides the qualitative feedback from participants, we adopted Spearman correlation (23) to study the relationship between the reported Influence on Valence (from end-of-day survey) and the average perceived valence for that day (from ESM responses). In this way, we can determine whether the lockdown had a positive or negative effect on valence. Likewise, we went through the same process in terms of users’ arousal and social roles. For social roles, we correlate the reported Influence on wlb (‘wlf’ indicates work-life balance) to the percentage of the day spent on working (reported in ‘work’ or ‘both’ states). For example, if someone claimed that the lockdown significantly influenced their work-life balance (i.e., a score of 5, which becomes 3 after Likert transformation in the pre-processing), and they spent 50% of the day at work, then (50%, 3) would be a point in the correlation.
5. Understanding App Usage Behaviors During COVID-19 Home confinement
@@ 5.2. Unsupervised Activity Recognition From App Sequences
Clustering. Next, AIGTS was conducted on all time series, resulting in a segmented list for each user. To extract features from these segments, k-means clustering (47) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette analysis (48) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
@@ 5.3. App Usage Correlation
Furthermore, we investigate the correlations between user activities with their valence, arousal, productivity, and interruptions. To understand whether any particular activity $A$ was correlated with valence, we ran a Spearman correlation (23) between valence scores and the amount of time spent in user activity $A$ in the previous hour. The same goes for the arousal, productivity, and interruptions, but the productivity and interruptions vary in duration from 1 hour to 24 hours, as these are calculated from the end-of-day survey and correspond to the value for a full day. Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day. The top correlation results for how different activities relate to usesr are displayed in Table 3. We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
As shown in Table 3, the correlations are weak with most activities, but few activities have moderate correlations. Tools related activities were positively correlated with overall valence and for both age groups. Particularly, Idle Tools have a relatively stronger correlation with valence in older adults, indicating an increase in the mood at work (Idle Tools include phone home screens and phone keyboards). Surfing by women and young age group are negatively correlated to valence, indicating reading news or surfing (e.g., reading COVID-19-related news) are negatively related to their valence. Productivity was negatively correlated with Communication Tools (e.g., keyboards and chat clients), and positively correlated with Business Communication (i.e., both Communication and Business apps such as Microsoft Teams). It can be seen that while both are Communication applications, whether or not they are used for work may imply productivity levels. Interruptions can be seen as the inverse of productivity, as it has positive correlation with Communication Tools, but the value is higher. The correlation reaches a moderate level for older adults while weakening for the young age group. This might cause by the difference in working loads between young and older age groups. Lifestyle negatively correlates with interruptions, especially for female and older adults.
@@
Table 4. The computed features from smartphone usage behaviors
@@ 6.1. Feature Extraction
Table 4 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (46; 76; 91), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transforme them into sine and cosine values. Additionally, it is worth noting that app-usage patterns exhibit variations between weekdays and weekends (40; 46), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notification is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app-usages, we employed our proposed method, AIGTS, as detailed in Section 5.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combing two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combing the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.

### v2

@@ 3.1. Background
To prevent the community spread of the COVID-19, Melbourne has been under six lockdowns since March 2020, with a total of 263 days of lockdowns, making it the city with the longest COVID-19 lockdown in the world (58; 85). Our data collection was conducted during the second lockdown period with Stage IV restrictions (the strictest measures), which lasted from 9 July 2020 to 27 October 2020, with our study conducted from 24 August 2020 to 13 September 2020 (see Figure 1). Among the six lockdown periods, this was one of the longest and strictest lockdown periods, during which a state of disaster was declared on 2 August 2020, followed by Stage IV restrictions in metropolitan Melbourne (62).
Under the Stage IV restrictions in force at the time of our data collection, a daily curfew was implemented from 8 PM to 5 AM, forbidding anyone from leaving their homes except for special reasons. Residents were only allowed to shop and exercise within 5 km of their homes, and all students returned to home-based learning. Other restrictions that were previously applied only to specific districts were extended to the entire Melbourne metropolitan area. On 13 September 2020, several restrictions were eased in Melbourne, including reduced curfews and loosened rules on outdoor exercise and social interactions. During home confinement due to the lockdown, people showed a strong reliance on technology, particularly for school and communication (74; 82). This presents a unique opportunity for researchers to gain insights into the psychological states and smartphone usage behaviours of individuals during home confinement. The knowledge gained could prove invaluable in understanding individuals in similar home confinement situations, such as those who are ill or undergoing post-recovery home rehabilitation, and home-based work.
@@ 3.2. Participants
Participants were recruited through advertising and comprised of university students, professional workers, academic staff, and health workers. In total, we recruited 32 participants (17 males and 15 females), where most participants were between 18-40 years old (mean = 25.44, median = 23.00, std = 8.95). The age distribution of participants across genders can be found in Figure 2. It is worth noting that participants were primarily recruited through advertisements on social media platforms Facebook and Discord, which may have resulted in an over-representation of users of these platforms in our sample. While this could be considered a limitation of our study, it is important to note that our research did not specifically focus on social media usage behaviour. Rather, we focus on individuals’ psychological states and smartphone usage behaviours during COVID-19 home confinement. Therefore, we believe that the potential oversample threat to the validity of our results is minimized. However, it is still important to note that our findings are specific to this particular population subset and may not be necessarily generalizable to other populations. Future research should consider recruiting participants from a wider range of sources to improve the generalizability of the findings.
After the data collection, each participant was compensated with a 25 AUD gift card as a token of appreciation for their time. Of the 32 participants, 27 recorded smartphone usage, 13 recorded desktop usage, and 8 recorded both smartphone and desktop app usage. Due to the relatively small number of participants using desktops, we primarily focused on analyzing the smartphone usage of the 27 participants. Analysis of desktop usage behaviours will be explored further in future work.
@@ 3.3. Procedures
Prior to data collection, participants were instructed to install the Balance for Android and Balance for Desktop applications (36) on their smartphones and desktops, respectively. These applications were specifically designed to record usage behaviours and enable background sensing and experience sampling, drawing inspiration from Van et al. (90). Both applications were optimized to minimize their consumption of computational power, energy, and storage resources. The applications utilized ESM, popping up a short questionnaire to participants every 90 minutes when their devices were actively in use between 7 am and 10 pm. This helps to obtain unbiased samples of participants’ activities, which served as the ground truth for our predictive models. Furthermore, an event-based ESM was triggered on the user’s smartphone after a minimum interaction of 10 minutes, with subsequent event-based ESMs occurring every 30 minutes.
@@ 3.3.1. Smartphone usage data
During the data collection, the background services of smartphones and desktops continuously tracked the following information: application usage (using Android Accessibility Services  [^2] Accessibility Services: https://developer.android.com/reference/android/accessibilityservice/AccessibilityService), notifications (using Android Notification Listeners  [^3] Notification Listeners: https://developer.android.com/reference/android/service/notification/NotificationListenerService), as well as the phone’s states (i.e. screen/battery status, ringer modes, location updates and physical activities). Specifically, the mobile app usage behaviour was recorded every time the user opened the app. For each notification, the application recorded the arrival time, hashed content (used for calculating the length of notification), the sender (private, work only, both or none), and participants’ relationship to the sender (family, friend, work and none). Table 1 displays an overview of the collected data for smartphone usage, ESM and end-of-day survey. In total, we have collected 502,485 recordings for smartphone usage data from 27 participants.
@@ 3.3.2. ESM responses
There are some popular questionnaires to measure well-being (43; 10), and the ESM questionnaire adapted from the Self-Assessment Manikin (SAM) (10) was employed in this research. In the ESM questionnaire, we collected information on valence, arousal, social role, interruptibility, and task from users. Especially, the item ‘valence’ was rated with a 5-point Likert scale from ‘unhappy’ to ‘happy’, and the item ‘arousal’ was also rated with a 5-point Likert scale from ‘calm’ to ‘excited’. Both valence and arousal were used to represent the users’ emotions based on the circumplex model (79). Understanding the social role, interruptiblity and task help us understand how people manage their time and attention on the digital devices (3). The detailed items of the ESM can be found in Table 1 and the screenshot of the ESM on users’ smartphones was shown in Figure 3(a). Overall, we have collected 1,749 ESM responses from 29 participants. Of these, 1,522 responses were obtained from 25 participants using smartphones (2 participants did not contribute any ESM responses), while the remaining 227 responses were from 10 participants using desktops.
@@ 4. Understanding Emotions, Social Roles, and Work-life-balance During COVID-19 Home Confinement
In this section, we aim to answer RQ1 ”What are people’s emotions and social roles like during COVID-19 home confinement? To what extent does the COVID-19 lockdown influence people’s emotions and work-life balance in their daily lives?”. To achieve this, we analyse 1522 ESM responses collected through smartphones from 25 participants, as described in Section 4.1. Furthermore, we explore the influence of the COVID-19 lockdown by examining the end-of-day survey in Section 4.2.
@@ 4.1. The Perceived Valence, Arousal and Social roles
The distribution of perceived valence, arousal, and social roles from the ESM responses is depicted in Figure 4. On average, participants reported a valence level of 3.46 (SD = 1.01) and an arousal level of 2.72 (SD = 1.11). Usually, the relatively high valance and low arousal refer to a state of calmness, relaxation and peace. This could be attributed to various factors during the COVID-19 home confinement. One possible reason is that people may feel relieved to be safe at home, and cherish the opportunity to spend more quality time with their family and loved ones, away from the potential risks of infection outside. Furthermore, the absence of the usual stressors of daily life, such as commuting, work pressures, and social obligations, may be contributing to a general sense of calm and well-being. Additionally, some people may have found comfort in the sense of community and solidarity that emerged during the pandemic, as people came together to support one another through difficult times. All of the factors mentioned earlier may contribute to the high valence and low arousal state during COVID-19 home confinement. Nevertheless, it is crucial to acknowledge that our data collection was limited to a three-week period, as COVID-19 Stage IV restrictions (strict home confinement) were lifted and steps for reopening commenced. Thereafter, in the event that strict home confinement lasts longer, it is possible that the emotional experiences of individuals may evolve as they adapt to the new normal and encounter new challenges. For instance, prolonged periods of confinement may lead to feelings of boredom, loneliness, and frustration.
@@
Next, we investigate the distribution of social roles and their correlation with emotion during home confinement. A social role refers to a set of norms, expectations, rules, beliefs and behaviours that are associated with a particular social situation (75). These roles can significantly influence users’ behaviours and the type of information they prefer to receive (3). As shown in Figure 4(b), smartphone users predominantly engaged in a private-related role during home confinement, which suggests that individuals tend to prioritize their personal lives and relationships while staying at home.
Since social roles play a crucial role in shaping individuals’ well-being, we investigate the emotions of users in different social roles, as shown in Figure 4(c). We found that, when engaging in private-related roles, participants had the highest levels of valence and lowest levels of arousal compared to work-related roles and hybrid roles. This may be because private-related roles involve personal activities which can provide a sense of comfort and happiness, while work-related roles or hybrid roles may involve stress, deadlines, and pressure to perform. The results of Welch’s t-tests (99) further support these findings, indicating that there is a significant difference in valence levels between private-related roles and the other two roles (i.e., private vs. both, private vs. work), with p-value less than 0.05. The above findings provide insights into how individuals can structure their daily routines to promote positive emotions during periods of home confinement, and interventions aimed at improving well-being during home confinement may benefit from considering the types of social roles that individuals are engaging in (34).
@@ 4.2. The Influence of the COVID-19 Lockdown on Emotions, Social Roles and Work-life Balance
In total, 54.3%, 51.3% and 54.7% of end-of-day responses displayed that the lockdown has influenced their valence, energy, and work-life balance to some extent, as shown in Figure 5. For the influence of lockdown on valence, 13 participants reported that it had a negative effect, claiming they ‘felt quite anxious and sad’ as ‘staying at home has made me a little sad’. They complained that they were unable to see friends (9 responses), see families (5 responses), go outside (10 responses), and feel bored (6 responses), e.g., ‘I thought I might be able to see friends, but I did not’, ‘I want to go outside, to the park’, ‘I’m missing my family’, ‘It’s hard to say how it’s specifically changed my happiness today because it’s the build-up of it has gone on for so long. I’m bored and sick of my housemates and it feels like time has stopped ’. Nevertheless, two felt the lockdown had a positive effect, claiming that ‘I am a lot happier’ and ‘I was able to maximize my time to do some exercise and transition to work immediately’.
@@
5. Understanding App Usage Behaviours During COVID-19 Home confinement
@@ 5.2. Unsupervised Activity Recognition From App Sequences
Clustering. Next, AIGTS was conducted on all-time series, resulting in a segmented list for each user. To extract features from these segments, k-means clustering (51) was applied. The distribution of different categories appearing in a segment was treated as an $n$ dimensional vector, where $n$ is the number of app categories. To identify the optimal number of segments, Silhouette analysis (52) was utilized with a maximum of 20 clusters. Especially, the label indicating which cluster a segment belongs to was used as a feature representing the activity within that segment. These segment labels were derived manually from the average distribution of apps in that cluster. In this paper, a segment that belongs to a particular cluster will be referred to as an ‘activity’.
@@ 5.3. App Usage Correlation
Furthermore, we investigate the correlations between user activities with their valence, arousal, productivity, and interruptions. To understand whether any particular activity $A$ was correlated with valence, we ran a Spearman correlation (24) between valence scores and the amount of time spent in user activity $A$ in the previous hour. The same goes for the arousal, productivity, and interruptions, but the productivity and interruptions vary in duration from 1 hour to 24 hours, as these are calculated from the end-of-day survey and correspond to the value for a full day. Although the ESM responses also measure interruptibility, we used the interruptibility item from the end-of-day survey to assess the intensity of how the COVID lockdown affected well-being, as it directly captures the interruptions experienced throughout the day. The top correlation results for how different activities relate to users are displayed in Table 3. We also report the top predicting activities for each score and demographic group (i.e., male, female, ages 18-24, ages 25+).
As shown in Table 3, the correlations are weak with most activities, but few activities have moderate correlations. Tools related activities were positively correlated with overall valence and for both age groups. Particularly, Idle Tools have a relatively stronger correlation with valence in older adults, indicating an increase in the mood at work (Idle Tools include phone home screens and phone keyboards). Surfing by women and young age group are negatively correlated to valence, indicating reading news or surfing (e.g., reading COVID-19-related news) are negatively related to their valence. Productivity was negatively correlated with Communication Tools (e.g., keyboards and chat clients), and positively correlated with Business Communication (i.e., both Communication and Business apps such as Microsoft Teams). It can be seen that while both are Communication applications, whether or not they are used for work may imply productivity levels. Interruptions can be seen as the inverse of productivity, as it has positive correlation with Communication Tools, but the value is higher. The correlation reaches a moderate level for older adults while weakening for the young age group. This might be caused by the difference in working loads between young and older age groups. Lifestyle negatively correlates with interruptions, especially for female and older adults.
@@
Table 4. The computed features from smartphone usage behaviours
@@ 6.1. Feature Extraction
Table 4 provides an overview of the features utilized in our research. Similar to previous mobile sensing studies (50; 81; 96), we extract contextual information, notifications-related information, physical activity, app, and activities inferred from apps. To process the app-usage data, we divide it into segments based on the ESM recording frequency (set at 90 minutes) and stored it in temporal sequences. Subsequently, we extract the features from these 90-minute segments. In order to handle cyclical features such as days (e.g., Monday) and times (e.g., Morning), we transform them into sine and cosine values. Additionally, it is worth noting that app usage patterns exhibit variations between weekdays and weekends (44; 50), prompting us to conduct separate analyses for these two categories. Regarding the time period per day, we divide it into six 4-hour segments, i.e., Early Morning, Morning, Afternoon, Evening, Night, and Late Night. Certain frequency-based features, such as notification-related or physical-activity-related features, are converted using quartiles, specifically the minimum, 25th percentile, 50th percentile, and 75th percentile. For instance, the number of interacted notifications is categorized into groups as ‘0’, ‘1-3’, ‘4-9’, and ‘10+’. Furthermore, the physical activities considered in our study only include active physical activities like bicycling, walking, and running.
To capture user activities related to app usage, we employed our proposed method, AIGTS, as detailed in Section 5.2. AIGTS allows us to extract both activities and their corresponding durations. Based on the extracted information, we compute the feature ‘entropy_user_activity’ by determining the proportion of each activity’s time portion within the entire 90-minute time interval. In addition to extracting individual features such as ‘num_noti’, we also generate composite features by combining two existing features. For example, the feature ‘num_communication’ is calculated by selecting and combining the notification sources. This composite feature counts the number of notifications originating the Business, Communication, or Social app categories. Notifications that do not belong to these app categories are counted as ‘app_noti’.